The average mass of 4 boys is 36.5kg.When one boy left,the average mass of the remaining boys is 36kg.What is the mass of the boy who has left?
1
Expert's answer
2011-10-13T13:16:21-0400
Let m1,m2,m3,m4 be the mass of first, second, third and forth boy respectively. Then,
(m1+m2+m3+m4)/4 = 36.5& (*)
(m1+m2+m3)/3 = 36 (**)
As we know from (**),
m1+m2+m3 = 36*3 = 108. Let's substitute is into (*):
(108+m4)/4 = 36.5
Hence, m4 = 36.5*4 - 108 = 38 kg.
So, the mass of the boy who has left is 38 kg.